Solve by using the substution method. Leave answers in reduced improper fraction form when necessary.
7x-4y =14
y =x-1
x= ?, y= ?

Answer by mananth(16949)

7x-4y =14
y =x-1
x= ?, y= ?
7x-4y =14
7x-4(x-1)=14
7x-4x+4=14
3x=14-4
x= 10/3
Ananth